OpenStudy (konradzuse):

Solve the differential equation dy/dx = 5-y?

OpenStudy (konradzuse):

I know if it was x I would just integrate, but with this y do I just send it over to the other side?

OpenStudy (anonymous):

that's a separable DE.... dy/dx = 5-y ---> \(\large \frac{dy}{5-y}=dx \) now just integrate both sides...

OpenStudy (konradzuse):

u = 5-y du = -1 so it could be -integral of u'/u = dx

OpenStudy (konradzuse):

but they want in terms of y soo.... hmm

OpenStudy (anonymous):

I will give you some example: dy/dx=5y with initial condition y(0)=2? dy/dx = 5y INT (1/y) dy = INT 5 dx ln(y) = 5x + c y = e^(5x + c) = a*e^(5x) y(0) = 2, 2 = a*e^0 = a y = 2e^(5x)

OpenStudy (anonymous):

but if you multiply and divide the y term for (-1) in the numerator you have the derivative of the denominator so you can put as -Ln(5-y)=x+C you take exponentials at both sides...